Find the slope of a line perpendicular to the line through the points (2,4) and (-12,8)
Answer by khurshidi(7)

Find the slope of a line perpendicular to the line through the points (2,4) and (-12,8)
Solution:
Two points given are
P = (2, 4) = (x1, y1)
Q = (-12, 8) = (x2, y2)
Slope of the line PQ
m1 = y2-y1/x2-x1 = 8 - 4/-12 - 2 = 4/-14 = -2/7
There fore
m1 = -2/7
The slope of the line perpendicular to the given line PQ will be the negative reciprocal of the slope of the line PQ
i.e.
m2 = -1/m1
(m2 is the slope of the line perpendicular to the slope of the given line PQ)
m2 = -1/-2/7
or
m2 = 7/2
This is the required slope of the line perpendicular to the line through the points (2, 4) and (-12, 8).
Hope the answer your question
Regards,
Khurshidi
http://sscom.netai.net
facebook.com/SSCOMMathemagics